OpenStudy (anonymous):

If sin(−x)= -.2, cos(−x)= -.5 and tan(−x)= 4.7, what does sec(−x)csc(−x)cot(−x)sin(−x) equal to?

OpenStudy (anonymous):

i got -5 but the computer says that this is incorrect so i'm not sure where i went wrong

OpenStudy (noelgreco):

Unless I'm missing something \[\frac{ \sin(-x) }{ \cos(-x) }= \tan(-x)\] which does not comport with your numbers

OpenStudy (noelgreco):

What the question supposes (I guess) is that you take the reciprocals of the given values, and multiply it out which gives 2.13

OpenStudy (anonymous):

i forgot to include that the question starts off by saying that Suppose: sin(x)=0.2 cos(x)=−0.5 tan(x)=−4.7

OpenStudy (anonymous):

@NoelGreco

OpenStudy (noelgreco):

Eating dinner - back shortly, but that's irrelevant. Take the reciprocals and you still get 2.13